A page for a journal contains 80 square inches of type written words. The height of a page is twice the width. If the Margin around the type is to be 2 inches uniformly, what are the dimensions of a page? Please include work. This will help me to break down the lesson and complete the remainder of my work. Thanks so much!!!
Answer by edjones(8007)

w=width of the page, L=Length of the page.
L=2w
width of print = w-4 because of margins.
Length of print = 2w-4
.
(w-4)(2w-4)=80
2w^2-12w+16=80
w^2-6w+8=40 divide each side by 2.
w^2-6w-32=0
w=9.4"
L=18.8"
.
Ed